write the equation of the line that contains the point (-3,4) and is perpendicular to the line y=-3x+1
Found 2 solutions by Greenfinch, ewatrrr:
Answer by Greenfinch(383)

Gradient of first line is -3, so line perpendicular is gradient of +1/3.
So line is 3y = x + c. Inserting point -3,4 gives 3x4 = -3 + c
So c = 15
Line is therefore 3y = x + 3

Answer by ewatrrr(24785)


Hi,
the standard point-slope form for an equation of a line y = mx + b
where m is the slope and b the y-intercept
perpendicular to the line y = -3x+1: that slope will be (1/3)
y = (1/3)x + b Using ordered pair(-3,4) to solve for b
4 = (1/3)*-3 + b
4 + 1 = b
5 = b
y = (1/3)x + 5